# Conceptualization, Methodology, Investigation, Data curation, Formal analysis, Writing - Original draft preparation

### v4

@@ 4 Discussion
A multi-layered review strategy is suggested to apply LLMs to security code review. As the results of RQ1 indicate, the existing general-purpose LLMs still fall significantly short of reaching the effectiveness of manual security code review. However, LLMs can serve as auxiliary tools to assist reviewers. The findings of RQ3 suggest that LLMs are more adept at identifying relatively simple security defects in shorter, logic-centric code written by developers with less project experience. We recommend a multi-layered review strategy to fully leverage the capability of LLMs - reviewers can first use LLMs to conduct an initial automatic review of those short and simple commits for security defects, then review both the submitted code and the LLM’s detection results to improve efficiency. This would allow reviewers to concentrate on an in-depth review of those complex commits, including lengthy code written by more experienced developers, thereby reducing the risk of missing sophisticated security defects. Additionally, where LLMs currently show limitations, these complex commits should be more heavily incorporated into datasets used for fine-tuning LLMs to improve their effectiveness in security code review.
@@
Providing a task scope is crucial for enhancing the performance of LLMs. Our findings indicate that incorporating a CWE list into prompts significantly improves the performance of various LLMs in security defect detection, which aligns with the results of Steenhoek et al. (2024). Providing a CWE list can effectively narrow the problem domain, guiding LLMs to focus on key areas, thereby increasing detection accuracy. However, in real-world security reviews, although with experience and domain knowledge, the reviewers lack prior knowledge of all potential types of security defects that may be present in the code. Therefore, constructing a comprehensive range of CWEs covering commonly encountered security defect types is necessary. It should also be noted that an overly exhaustive scope can diminish the accuracy of LLM responses. In Bakhshandeh et al. (2023), providing a list of all 75 CWEs in their prompts decreased the accuracy of security defect detection and location by the LLMs. This drop in accuracy may be attributed to the overly lengthy and detailed list of CWEs provided, which interferes with the LLM’s capability to focus effectively on the target defect, yielding adverse effects. Due to the limited context length of the used LLMs, inputting the entire codebase during prompt design was not feasible. The lack of code context made it difficult for the LLM to fully understand the dependencies and interaction logic between files, which in turn affected its performance in identifying security defects, such as Race Condition and Insecure Library Version. We tested various approaches, including one-shot and few-shot prompting strategies, as well as providing the complete patchset within the prompt. These strategies required splitting the code into sections for input, which we found significantly degraded the performance of LLMs compared to the basic prompt $\textbf{P}_{b}$. In order to provide as much effective information as possible within the limited context window, we designed prompts that combined multiple elements, including project metadata, CWEs, code diffs, and the commit messages. We also standardized the output format to help the model better understand the task and introduced CoT reasoning to guide the model in inferring missing context. Ultimately, we identified and selected four prompts that outperformed others, which, together with the basic prompt, were selected as templates for our formal experiments. The results demonstrate that including a list of CWEs significantly enhances LLM performance, making it our study’s most effective prompting strategy. As the context window of LLMs continues to expand and prompt engineering techniques evolve, we plan to explore additional strategies in the future, such as incorporating RAG for external information and employing automated prompt optimization to dynamically adjust prompts during reasoning, thereby further improving the capability of LLMs in security defect detection.
Using LLMs in security code review requires developers to standardize their code commenting practices. According to RQ3, annotations containing information related to non-target security defects, which were often already resolved by developers, negatively impact the LLM’s capability in security code review. This suggests that LLMs could mistakenly interpret these resolved security defects as existing defects. Moreover, although the p-value of ‘AnnotationRatio’ (0.21) is more than 0.05, we still consider its OR value (0.56) to be somewhat informative. The more annotations in the code, the more challenging it is for LLMs to detect security defects, reflecting the disruptive effect and poor quality of annotations written by developers. Hence, we recommend that developers write code comments in a standardized way to clarify the purpose of each comment, thus better supporting LLMs in code review tasks.
@@
With the development of LLMs, while expanding its context window, it is also necessary to ensure that LLMs can accurately capture details from long inputs. The emergence of LLMs that support large amounts of tokens has alleviated the challenge of inputting code context information in applying LLMs to code-related tasks. However, as evidenced by the results of RQ3, when dealing with code that contains more tokens, it is more difficult for LLMs to identify target security defects. This is because many security defects are introduced by small code snippets, and longer inputs may distract LLMs from overlooking certain details, thereby failing to detect security defects introduced by small code snippets. Additionally, according to RQ1, despite GPT-4 Turbo having a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and Llama 2. Therefore, ensuring the accuracy of security defect detection when expanding the LLMs’ context window is a promising direction.
@@
A key challenge in applying LLMs to security code review is ensuring comprehensive detection of security defects while maintaining consistent results. The results from RQ1.4 indicate that LLMs with stronger capabilities to detect security defects tend to exhibit poorer consistency in their responses. This may be because the outputs generated by LLMs with higher non-determinism are more diverse, making them more likely to include content related to target security defects compared to those generated by LLMs with lower non-determinism. Both detection capability and the stability of detection results are crucial for applying LLMs in security code review. However, current studies related to LLMs in vulnerability detection have largely overlooked the importance of the stability of results. Therefore, for researchers of LLMs, exploring approaches to balance detection capabilities with the non-determinism inherent in LLMs represents a valuable direction for future improvements.
@@ 5 Threats to Validity
Internal Validity refers to whether an observed association between two variables can be attributed to a causal link between them. One of the internal validity threats could be data leakage. The data used in our study was obtained from two large-scale open-source projects, which may overlap with the data used for training LLMs, a common problem with LLM-based studies. Since the training data cutoff dates for the models used (GPT-3.5 and GPT-4) are not disclosed and the dataset we used  (Yu et al., 2023) predates the release of some of the LLMs we selected, we cannot avoid the overlap between the dataset we used and the training data of LLMs by filtering the data for a specific period of time. Moreover, according to Cao et al. (2024), contaminated data does not always affect the results, and the model sometimes even performs better after the model’s cutoff date. Another threat is the design of prompts. We design our prompt according to best practices to evoke the best model performance. We selected words that are as unbiased, non-suggestive, and unambiguous as possible. We iteratively revised the prompts to avoid individual words from impacting the responses of LLMs during our preliminary exploration. In addition, we only included code from the single vulnerable file rather than the whole patchset in the prompt, which may affect the performance of LLMs due to the lack of contextual information. To specifically analyze its impact, we designed a prompt $\textbf{P}_{cot}$, in which LLMs are instructed to generate code context according to the commit message before detecting security defects.
@@
External Validity concerns whether a causal link generalizes across contexts. In this study, the experiments were conducted on 6 LLMs based on 4 open-source projects across programming languages including Python and C/C++. We acknowledge that the dataset we utilized may not fully represent industrial projects across various programming languages, and the performance of the selected LLMs may not represent the capabilities of all existing LLMs. We implemented the following strategies to ensure the generalizability of the experimental results. For model diversity, our experiments incorporated both open-source and proprietary LLMs of varying parameter scales to encompass a diverse range of popular LLMs. For dataset construction, we collected code review data from real-world open-source projects in the OpenStack and Qt communities. These projects differ in architectures, programming languages, and community structures, providing a comprehensive and valuable dataset for empirical analysis.
@@
Construct Validity assesses whether a specific set of metrics corresponds to what they are intended to measure. To measure the performance of LLMs, we classified the natural language responses of LLMs into four predefined categories, with human reviews of security defects in code as the ground truth. There is a potential natural language understanding bias which affects classification accuracy. To reduce the bias, we established granular definitions for each category in three dimensions: defect type, location, and description, supplemented by examples for each category to ensure that the metrics are accurate, appropriate, and consistent throughout the evaluation process. In addition, the data labeling in RQ1 and the analysis of quality problems in RQ2 were all carried out manually. In order to further reduce the impact of subjective judgments, we employed open coding and constant comparison (Glaser, 1965). In the data labeling employed for RQ1, the first and third authors conducted a pilot data labeling and completed all the labeling work after reaching a consensus on the criteria through discussion. In the data analysis of RQ2, the first and seventh authors, conducted two rounds of pilot data analysis to determine the definition of each quality problem and achieved a high level of consistency in the results. All discrepancies in the manual work of our study were reviewed and resolved through negotiation and discussion among at least three authors, thus ensuring the reliability of the findings.
@@ 6.1 Security Defect Detection in Code Review
Several studies have analyzed security defect detection during code reviews. Paul et al. (2021) built a dataset from the Chromium OS project, consisting of 516 code reviews that successfully identified security defects and 374 code reviews that missed security defects. The study analyzed 18 code review attributes, including review duration, reviewer experience, and author experience, to explore factors influencing the identification of security defects. It was found that the duration of code reviews and the frequency of peer reviews positively affected the detection of security defects, while the number of directories under review and the number of prior commits had a negative impact. Alfadel et al. (2023) examined pull requests (PR) from 10 popular JavaScript GitHub projects within the npm ecosystem, then analyzed the frequency, categories, and handling of security defects identified during code reviews. Their work indicated that only a small fraction of PRs raise security problems, and most of the identified security problems were resolved during the code review.
@@ 6.2 Code Characteristics of Security Defects
Alves et al. (2016) constructed a security defect dataset based on five widely used projects and 27 software metrics, such as lines of code, cyclomatic complexity and knots in the control flow graph, among others. This study demonstrated that most of these metrics are effective in distinguishing vulnerable from non-vulnerable code, thereby providing valuable support for security practitioners. Meneely et al. (2013) conducted quantitative and qualitative analyses on 124 Vulnerability-Contributing Commits (VCC) in the Apache HTTP Server to identify characteristics of code changes with security defects. The findings suggest that longer lines of code and less experienced committers are more likely to introduce security defects. Bosu et al. (2014) identified 413 VCCs from code reviews in 10 open-source projects and reported consistent observations with Meneely et al. (2013) that longer lines of code and less experienced committers are more likely to introduce security defects. Additionally, they indicated that modified code files are more likely to contain security defects than new files.
@@ 6.3 Automated Tools of Security Defect Detection
Charoenwet et al. (2024) conducted an empirical study on the application of static application security testing tools (SASTs) in security code review and found that a single SAST tool could produce warnings for 52% of VCCs. By combining several tools, warnings could be produced for 78% of the VCCs. However, at least 76% of these warnings are irrelevant to the vulnerability in VCCs, highlighting the challenges that SAST tools face in providing detailed information about the vulnerability during security code review. Inspired by this work, our work investigated the application of LLMs in security code review, aiming to overcome the limitations of SAST tools to some extent.
@@
Several studies have explored the performance of LLMs in detecting security defects. The majority of these studies conducted a coarse-grained assessment of LLMs in binary classification tasks that determine whether the code contains any security defects without specifying the corresponding defect type. Zhou et al. (2024b) utilized Common Weakness Enumeration (CWE), project information, and a few-shot learning approach to design prompts and evaluated the capability of two popular LLMs, GPT-3.5 and GPT-4, in binary judgment under various prompts. They found that GPT-3.5 achieved competitive performance with the prior state-of-the-art detection approaches, while GPT-4 consistently outperformed the state-of-the-art. Purba et al. (2023) compared the performance of LLMs with popular static analysis tools in binary detection. Their work demonstrates that although LLMs achieved a good recall rate, their false positive rate was significantly higher than that of static analysis tools. Unlike previous studies, our work provides a more fine-grained evaluation of the capability of LLMs in security code reviews by including additional insights such as defect localization, description, and suggestions for defect remediation.
@@
Previous studies also investigated whether LLMs could provide more information in security defect detection. Bakhshandeh et al. (2023) provided LLMs with a list of CWEs and the inspection results of static analysis tools on the prompts, requesting the LLMs to generate specific CWE names along with the corresponding code line numbers. Their study shows that LLMs effectively reduce false positives of static tools and demonstrate competitive or even superior performance in defect localization compared to these tools. Yin and Ni (2024) evaluated the performance of LLMs across multiple tasks in security analysis, including defect detection, localization, and description, and found that the existing state-of-the-art approaches, such as LineVul (Fu and Tantithamthavorn, 2022) and SVulD (Ni et al., 2023), generally outperformed LLMs. Specifically, LLMs exhibited limited overall accuracy and varied performance across different CWE types in defect localization and description tasks. These studies have all been limited to quantitatively evaluating LLM performance. In comparison, our work performed an in-depth analysis of the quality problems in LLM-generated responses and factors affecting LLM performance to understand the potential and challenges LLMs face in security code review.
Employing specific prompting frameworks to enhance the performance of LLMs in identifying security defects has also been explored. Yang et al. (2024) introduced a framework utilizing DL models, integrating context learning and Chain-of-Thought (CoT) to augment prompts in security defect detection. Their framework yielded superior performance compared to state-of-the-art prompting techniques. Nong et al. (2024a) proposed a vulnerability-semantics-guided prompting (VSP) approach that combines CoT with various auxiliary information to evaluate the performance of LLMs in identifying security defects (including their classification and resolution). The research concerns and employed datasets of the above two studies (Yang et al., 2024; Nong et al., 2024a) differ significantly from ours. These studies focus on designing benchmarks to enhance LLM in vulnerability detection, utilizing synthetic datasets and CVE datasets composed of publicly disclosed vulnerabilities to validate their improvements. Compared with them, our work focuses on the empirical evaluation of LLMs in the context of security code review, employing a dataset of security defects collected from code reviews of real-world software projects.
It is worth noting that the studies mentioned above focus on general vulnerability detection by LLMs, whereas our work targets the specific area of security code review by LLMs. Inspired by these related works, we aim to leverage LLMs’ potential in security analysis to identify the obstacles, application strategies, and improvement directions for integrating LLMs into security code review.

### v5

@@ 4 Discussion
A multi-layered review strategy is suggested to apply LLMs to security code review. As the results of RQ1 indicate, the existing general-purpose LLMs still fall significantly short of reaching the effectiveness of manual security code review. However, LLMs can serve as auxiliary tools to assist reviewers. We recommend a multi-layered review strategy to fully leverage the capability of LLMs. The findings of RQ3 suggest that LLMs are more adept at identifying security defects in code file with fewer tokens and higher cyclomatic complexity. Utilizing these two factors, a lightweight model can be trained to classify the code file to be reviewed into the low-difficulty or high-difficulty groups. Firstly, LLMs are utilized to conduct an initial automatic review of these files. Then, for low-difficulty code, reviewers quickly confirm the security defects identified by LLMs. Code in which LLMs did not detect any security defects would then be resampled and rechecked to confirm its security. For high-difficulty code, reviewers utilized the detection results of LLMs as a reference, combined with their own domain knowledge, to perform a more thorough and in-depth analysis of the submitted code. This approach aims to improve the efficiency of security code review while minimizing the risk of missing security defects. Additionally, reviewers should provide feedback on each of the detection results generated by LLMs, which can then be used to train both the difficulty classification model and LLMs for security code review.
@@
When using a CWE list as auxiliary information for LLM in security code review, the granularity of CWE entries is crucial. Our findings indicate that incorporating a CWE list into prompts significantly improves the performance of various LLMs in security defect detection, which aligns with the results of steenhoek2024comprehensive. However, in real-world security reviews, although with experience and domain knowledge, the reviewers lack prior knowledge of all potential types of security defects that may be present in the code. It should also be noted that an overly exhaustive scope can diminish the accuracy of LLM responses. In bakhshandeh2023using, providing a list of all 75 CWEs in their prompts decreased the accuracy of security defect detection and location by the LLMs. This drop in accuracy may be attributed to the overly lengthy and detailed list of CWEs provided, which interferes with the LLM’s capability to focus effectively on the target defect, yielding adverse effects. Therefore, constructing a concise yet comprehensive CWE list is necessary. In our experiments, due to the limited context length of the used LLMs, inputting the entire codebase during prompt design was not feasible. The lack of code context made it difficult for the LLM to fully understand the dependencies and interaction logic between files, which in turn affected its performance in identifying security defects, such as Race Condition and Insecure Library Version. We tested various approaches, including one-shot and few-shot prompting strategies, as well as providing the complete patchset within the prompt. These strategies required splitting the code into sections for input, which we found significantly degraded the performance of LLMs compared to the basic prompt $\textbf{P}_{b}$. In order to provide as much effective information as possible within the limited context window, we designed prompts that combined multiple elements, including project metadata, CWEs, code diffs, and the commit messages. We also standardized the output format to help the model better understand the task and introduced CoT reasoning to guide the model in inferring missing context. Ultimately, we identified and selected four prompts that outperformed others, which, together with the basic prompt, were selected as templates for our formal experiments. The results in RQ1.3 demonstrate that for GPT-4, including a CWE list in the prompt significantly enhances LLM performance, while for DeepSeek-R1, the impact of the provided CWE list on model performance depends on the specific programming language. We therefore emphasize the need to explore CWE entries of different granularities for various types of security defects to determine the optimal CWE list. As the context window of LLMs continues to expand and prompt engineering techniques evolve, we also plan to explore additional strategies in the future, such as incorporating RAG for external information, employing automated prompt optimization to dynamically adjust prompts, and leveraging the agent architecture to dynamically invoke contextual information during reasoning, thereby further improving the capability of LLMs in security defect detection.
@@
With the development of LLMs, while expanding its context window, it is also necessary to ensure that LLMs can accurately capture details from long inputs. The emergence of LLMs that support large amounts of tokens has alleviated the challenge of inputting code context information in applying LLMs to code-related tasks. However, as evidenced by the results of RQ3, when dealing with code that contains more tokens, it is more difficult for LLMs to identify target security defects. This is because many security defects are introduced by small code snippets, and longer inputs may distract LLMs from overlooking certain details, thereby failing to detect security defects introduced by small code snippets. Additionally, according to RQ1, for general-purpose LLMs, despite GPT-4 Turbo having a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and Llama 2. While for the reasoning-optimized LLM, although DeepSeek-R1 also has a 128k context window and is the best-performing model in this study, its rate of Incorrect fact reaches as high as 31.81%. Specifically, DeepSeek-R1 often presented problems with mismatched code line numbers and code snippets in its responses, which significantly undermines the understandability of its analysis. Therefore, while expanding the LLMs’ context window, it is also necessary to guarantee the model’s precise memory of detailed information such as key code segments, to enhance the accuracy of security defect detection and the clarity of defect descriptions.
@@
A key challenge in applying LLMs to security code review is ensuring comprehensive detection of security defects while maintaining consistent results. The results from RQ1.4 indicate that the two LLMs (DeepSeek-R1 and GPT-4) with stronger capabilities to detect security defects tend to exhibit poorer consistency in their responses. This may be because LLMs with better performance are more capable of reasoning from multiple perspectives. The inherent randomness of LLM leads to variations in reasoning paths, resulting in inconsistent outcomes. In comparison, the thought of less capable models is relatively shallow, and their strong consistency in responses may stem from failing to deeply analyze the code. The security code review task requires models to explore as many reasoning paths as possible in order to maximize the recall of security defects, making the inconsistency of responses inevitable. In real-world application of LLMs for security code review, both the accuracy and stability of detection results are crucial. However, current studies related to LLMs in vulnerability detection have largely overlooked the importance of the stability of results. Therefore, for researchers, exploring approaches to improve the detection sensitivity while maintaining result stability, such as through more precise contextual information or voting mechanisms, is a valuable direction for future improvement.
Reasoning models exhibit limited sensitivity to logic-related security defects during security code review The results from RQ3 show that DeepSeek-R1 underperformed in detecting race conditions compared to most other security defect types. Previous studies have also demonstrated that the detection accuracy of LLMs suffers from an imbalance across different security defect types, with race conditions being relatively underexplored sheng2025llms. Given that Race Condition is a high-frequency type of security defect in Python projects (accounting for 65.1% on the Python dataset in this study, see fig. 3), it is crucial to guide LLMs to pay more attention to logic-related security defects. Based on the results of RQ1, $\textbf{P}_{cid}$ and $\textbf{P}_{cot-guardrail}$ enhanced DeepSeek-R1’s capability of detecting security defect detection. To separately analyze their effect on race conditions, we further calculated the IH-Scores of DeepSeek-R1 on code files with race conditions from the complete dataset (Python and C/C++) under $\textbf{P}_{b}$, $\textbf{P}_{cid}$, and $\textbf{P}_{cot-guardrail}$, which were 5.09%, 7.56%, and 6.02%, respectively. This suggests two potentially effective approaches to improve the sensitivity of LLMs on logical-related security defects: first, explicitly guiding the model to concentrate on the corresponding areas through CWE entries; and second, leveraging the commit message to assist LLMs in better understanding code functionality and logic implementation.
@@ 5 Threats to Validity
Internal Validity refers to whether an observed association between two variables can be attributed to a causal link between them. One of the internal validity threats could be data leakage. The data used in our study was obtained from two large-scale open-source projects, which may overlap with the data used for training LLMs, a common problem with LLM-based studies. Since the training data cutoff dates for the models used (GPT-3.5 and GPT-4) are not disclosed and the dataset we used  (yu2023security) predates the release of some of the LLMs we selected, we cannot avoid the overlap between the dataset we used and the training data of LLMs by filtering the data for a specific period of time. Moreover, according to cao2024concerned, contaminated data does not always affect the results, and the model sometimes even performs better after the model’s cutoff date. Another threat is the design of prompts. We design our prompt according to best practices to evoke the best model performance. We selected words that are as unbiased, non-suggestive, and unambiguous as possible. We iteratively revised the prompts to avoid individual words from impacting the responses of LLMs during our preliminary exploration. In addition, we only included code from the single vulnerable file rather than the whole patchset in the prompt, which may affect the performance of LLMs due to the lack of contextual information. To explore approaches for context augmentation, we designed a prompt $\textbf{P}_{cot}$, in which LLMs are instructed to reason step-by-step—generating code context according to the commit message before detecting security defects. Since multiple factors in $\textbf{P}_{cot}$ could influence model performance, we also designed $\textbf{P}_{cot-guardrail}$ as a control prompt that includes only the commit message and CoT instruction, in order to isolate the effect of the fabricated code context.
@@
External Validity concerns whether a causal link generalizes across contexts. In this study, the experiments were conducted on 7 LLMs based on 4 open-source projects across programming languages including Python and C/C++. We acknowledge that the dataset we utilized may not fully represent industrial projects across various programming languages, and the performance of the selected LLMs may not represent the capabilities of all existing LLMs. We implemented the following strategies to ensure the generalizability of the experimental results. For model diversity, our experiments incorporated both open-source and proprietary LLMs of varying parameter scales to encompass a diverse range of popular general-purpose and reasoning-optimized LLMs. For dataset construction, we collected code review data from real-world open-source projects in the OpenStack and Qt communities. These projects differ in architectures, programming languages, and community structures, providing a comprehensive and valuable dataset for empirical analysis.
@@
Construct Validity assesses whether a specific set of metrics corresponds to what they are intended to measure. To measure the performance of LLMs, we classified the natural language responses of LLMs into four predefined categories, with human reviews of security defects in code as the ground truth. There is a potential natural language understanding bias which affects classification accuracy. To reduce the bias, we established granular definitions for each category in three dimensions: defect type, location, and description, supplemented by examples for each category to ensure that the metrics are accurate, appropriate, and consistent throughout the evaluation process. In addition, the data labeling in RQ1 and the analysis of quality problems in RQ2 were all carried out manually. In order to further reduce the impact of subjective judgments, we employed open coding and constant comparison (glaser1965constant). In the data labeling employed for RQ1, the first and third authors conducted a pilot data labeling and completed all the labeling work after reaching a consensus on the criteria through discussion. In the data analysis of RQ2, the first and seventh authors, conducted two rounds of pilot data analysis to determine the definition of each quality problem and achieved a high level of consistency in the results. All discrepancies in the manual work of our study were reviewed and resolved through negotiation and discussion among at least three authors, thus ensuring the reliability of the findings.
@@ 6.1 Security Defect Detection in Code Review
Several studies have analyzed security defect detection during code reviews. paul2021security built a dataset from the Chromium OS project, consisting of 516 code reviews that successfully identified security defects and 374 code reviews that missed security defects. The study analyzed 18 code review attributes, including review duration, reviewer experience, and author experience, to explore factors influencing the identification of security defects. It was found that the duration of code reviews and the frequency of peer reviews positively affected the detection of security defects, while the number of directories under review and the number of prior commits had a negative impact. alfadel2023empirical examined pull requests (PR) from 10 popular JavaScript GitHub projects within the npm ecosystem, then analyzed the frequency, categories, and handling of security defects identified during code reviews. Their work indicated that only a small fraction of PRs raise security problems, and most of the identified security problems were resolved during the code review.
@@ 6.2 Code Characteristics of Security Defects
alves2016software constructed a security defect dataset based on five widely used projects and 27 software metrics, such as lines of code, cyclomatic complexity and knots in the control flow graph, among others. This study demonstrated that most of these metrics are effective in distinguishing vulnerable from non-vulnerable code, thereby providing valuable support for security practitioners. meneely2013patch conducted quantitative and qualitative analyses on 124 Vulnerability-Contributing Commits (VCC) in the Apache HTTP Server to identify characteristics of code changes with security defects. The findings suggest that longer lines of code and less experienced committers are more likely to introduce security defects. bosu2014identifying identified 413 VCCs from code reviews in 10 open-source projects and reported consistent observations with meneely2013patch that longer lines of code and less experienced committers are more likely to introduce security defects. Additionally, they indicated that modified code files are more likely to contain security defects than new files.
@@ 6.3 Automated Tools of Security Defect Detection
charoenwet2024empirical conducted an empirical study on the application of static application security testing tools (SASTs) in security code review and found that a single SAST tool could produce warnings for 52% of VCCs. By combining several tools, warnings could be produced for 78% of the VCCs. However, at least 76% of these warnings are irrelevant to the vulnerability in VCCs, highlighting the challenges that SAST tools face in providing detailed information about the vulnerability during security code review. Inspired by this work, our work investigated the application of LLMs in security code review, aiming to overcome the limitations of SAST tools to some extent.
@@
Several studies have explored the performance of LLMs in detecting security defects. The majority of these studies conducted a coarse-grained assessment of LLMs in binary classification tasks that determine whether the code contains any security defects without specifying the corresponding defect type. zhou2024largelanguage utilized CWE, project information, and a few-shot learning approach to design prompts and evaluated the capability of two popular LLMs, GPT-3.5 and GPT-4, in binary judgment under various prompts. They found that GPT-3.5 achieved competitive performance with the prior state-of-the-art detection approaches, while GPT-4 consistently outperformed the state-of-the-art. purba2023software compared the performance of LLMs with popular static analysis tools in binary detection. Their work demonstrates that although LLMs achieved a good recall rate, their false positive rate was significantly higher than that of static analysis tools.  jensen2024software evaluated the performance of three proprietary OpenAI LLMs and four small-scale models with fewer than 13B parameters to identify vulnerabilities in the SecurityEval dataset (siddiq2022securityeval). They adopted zero-shot and CoT prompting and found that the best‑performing model - GPT‑4 - achieved an accuracy of 95.6% in distinguishing between vulnerable and patched versions of code. Similarly,  steenhoek2024err assessed the capabilities of 14 LLMs on a real-world and function-level vulnerability dataset, SVEN, across seven different prompt strategies. They explored various approaches, including prompt optimization, scaling model size, increasing pre-training data, fine-tuning, and incorporating domain knowledge into prompts, yet none led to significant improvements in the performance of models in the binary vulnerability detection task. Unlike previous studies, our work provides a more fine-grained evaluation of the capability of LLMs in security code reviews by including additional insights such as defect localization, description, and suggestions for defect remediation.
@@
Previous studies also investigated whether LLMs could provide more information in security defect detection. bakhshandeh2023using provided LLMs with a list of CWEs and the inspection results of static analysis tools on the prompts, requesting the LLMs to generate specific CWE names along with the corresponding code line numbers. Their study shows that LLMs effectively reduce false positives of static tools and demonstrate competitive or even superior performance in defect localization compared to these tools. yin2024multitask evaluated the performance of LLMs across multiple tasks in security analysis, including defect detection, localization, and description, and found that the existing state-of-the-art approaches, such as LineVul (fu2022linevul) and SVulD (ni2023distinguishing), generally outperformed LLMs. Specifically, LLMs exhibited limited overall accuracy and varied performance across different CWE types in defect localization and description tasks.  khare2025understanding comprehensively evaluated 16 pre-trained LLMs for detecting security defects on a function-level vulnerability dataset of 5,000 samples. Their study systematically compared model performance across three distinct prompting strategies and found that involving step-by-step analysis significantly improved LLM performance. This work also performed comparison of LLMs against leading static analysis tools and deep learning-based approaches, finding that LLMs reported higher detection accuracies than them. These studies have all been limited to quantitatively evaluating LLM performance. In comparison, our work performed an in-depth analysis of the quality problems in LLM-generated responses and factors affecting LLM performance to understand the potential and challenges LLMs face in security code review.
Fine-tuning has also been employed to improve the capability of LLMs in security defect detection.  mechri2025secureqwen fine-tuned CodeQwen1.5-7B on a large Python dataset comprising both real-world code and synthetic code generated by LLMs, in order to guide the model to capture and classify security defects in code accurately. The model ultimately achieved high accuracy, with F1 scores ranging from 84% to 99%. Although fine-tuning can effectively improve the accuracy of model, fine-tuned LLMs often exhibit poor generalization. In practical application, adjustments to the model parameters and subsequent retraining are usually required. Consequently, some studies have shifted their focus toward approaches that avoid altering the core model, such as leveraging prompt engineering or incorporating external information to further enhance the precision of LLMs in security defect detection. yang2024dlap introduced a framework utilizing DL models, integrating context learning and CoT to augment prompts in security defect detection. Their framework yielded superior performance compared to state-of-the-art prompting techniques. nong2024chain proposed a vulnerability-semantics-guided prompting (VSP) approach that combines CoT with various auxiliary information to evaluate the performance of LLMs in identifying security defects (including their classification and resolution). The research concerns and employed datasets of the above two studies (yang2024dlap; nong2024chain) differ significantly from ours. These studies focus on designing benchmarks to enhance LLM in vulnerability detection, utilizing synthetic datasets and CVE datasets composed of publicly disclosed vulnerabilities to validate their improvements. Compared with them, our work focuses on the empirical evaluation of LLMs in the context of security code review, employing a dataset of security defects collected from code reviews of real-world software projects. To further improve the accuracy of LLMs in security defect detection, some studies have adopted RAG to retrieve external knowledge and enhance the input prompts. du2024vul constructed multi-dimensional representations to retrieve knowledge most relevant to the target code from historical vulnerabilities and their corresponding fixes. This framework improved the accuracy of LLMs in distinguishing vulnerable code from patched code by 16%–24%. lu2024grace incorporated graph-structured information from the code and identified highly related code samples as external guidance into the prompt. Using this method, LLMs outperformed state-of-the-art deep learning-based tools in detecting vulnerabilities. With the emergency of reasoning-optimized LLMs, LLM-based agents have also shown promising applications in security defect detection. Inspired by courtroom, widyasari2025let constructed a framework consisting of four distinct agent roles, which are security researcher, code author, moderator, and review board. Their framework employs multiple rounds of interaction to arrive at a final determination on security defects in the code, which substantially surpassed the performance of approaches based on CoT and fine-tuning. tsai2025leveraging incorporated RAG into the multi-turn multi-role interaction agent framework, achieving a 29.9% improvement in F1 score compared to widyasari2025let. It is worth noting that the studies mentioned above focus on general vulnerability detection by LLMs, whereas our work targets the specific area of security code review by LLMs. Inspired by these related works, we aim to leverage LLMs’ potential in security analysis to identify the obstacles, application strategies, and improvement directions for integrating LLMs into security code review.